OpenStudy (anonymous):

If f(x) = 5x – 4 and g(x) = 6x + 3, what is f(g(x))?

OpenStudy (btaylor):

f(g(x) = 5(6x+3) - 4 = 30x+15 - 4 = 30x + 11